OpenStudy (anonymous):

Can you check and correct my answers for Algebra?! 1- 8n/8 2- 2x-2/10 3-b 4-a 5-c 6- 3x/4 7- 3p/6 8- x+11/7

OpenStudy (anonymous):

Can you post it with another format, because I cant open it.

OpenStudy (anonymous):

The first one is wrong, what did you do to get to this result?

OpenStudy (anonymous):

Is that \[\frac{8n}{8}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

would #1 be 11n/4

OpenStudy (anonymous):

So, when summing fractions, you think like a pizza, I know this sounds childish, but that is actually all you need. If you have a quarter of a pizza, and you add another one, you end up with two quarters:\[\frac{1}{4}+\frac{1}{4}=\frac{2}{4}\]And not:\[\frac{2}{8}\]So, when summing fractions, the bottom remais the same (and both need to have the same number on the bottom), and the upper part is summed, so on the first one you have:\[\frac{8n}{4}+\frac{3n}{4}=\frac{11n}{4}\]Did you understand that?

OpenStudy (anonymous):

YES thank you

OpenStudy (anonymous):

for number 2 would it be (2x - 2)/5

OpenStudy (anonymous):

Yeah, thats it

OpenStudy (anonymous):

What would #3be?

OpenStudy (anonymous):

Would it be choice B?

OpenStudy (anonymous):

This one is a little bit harder, its not choice b, after you sum, try to make it simpler, for example: 5x+10=5(x+2), but with different numbers.

OpenStudy (anonymous):

I've tried but thats the answer I got

OpenStudy (anonymous):

b is correct by the way, but is not the simplest form, wich is what he asks for.

OpenStudy (anonymous):

Oh okay, And #4 is C?

OpenStudy (anonymous):

Yeah, thats correct. Did you got to A on #3?

OpenStudy (anonymous):

No I got b

OpenStudy (anonymous):

I'm really stuck on #5 and #8

OpenStudy (anonymous):

Ok, we come back later to #3, #5 is correct, but since you said you were stuck I assume you didnt understand it. So, on a all positives and negatives are switched. You can look at it this way:\[\frac{4}{3-x}=1\left(\frac{4}{3-x}\right)\]But 1 can be written like\[\frac{-1}{-1}\]Thats what they do there, to change the signs. You see why you don't change anything when you change the signs on a division? The second one is just a different way to write the first and doesn't change anything. Now the third one, is wrong, do you see why?

OpenStudy (anonymous):

No I;m not sure why

OpenStudy (anonymous):

Ok, the first and the second are the same, because you multiplied the top part AND the bottom part by the same number, wich is the same as multiplying everything by 1. Now on the c choice, the bottom part continues the same, and only the upper part is multiplied, that makes it take the opposite value, the negative one, because you are not multiplying the bottom too.

OpenStudy (anonymous):

Oh ok i get it now! Thanks. For #7 would it be 1p/6

OpenStudy (anonymous):

Yep, thats the answer of #7

OpenStudy (anonymous):

What would be the answer for 8?

OpenStudy (anonymous):

Wait, we skipped #6

OpenStudy (anonymous):

6 is 5x/4

OpenStudy (anonymous):

Yep, now on the #8, you are subtracting *everything* on the right side, from everything on the left side. The operation is the same as in 7, but with more terms.

OpenStudy (anonymous):

I get that, but i just can't get the answr

OpenStudy (anonymous):

is it x+5/7

OpenStudy (anonymous):

So, the part where you are having trouble with is:\[(x+8)-(x-3)\] When you have something like this, eberything on the right side is multiplied by -1, so the same thing can be written like this:\[(x+8)+(-x+3)\]Do you see it now?

OpenStudy (anonymous):

So I would add the two together...

OpenStudy (anonymous):

For 9 would it be A and for 10 would it be C?

OpenStudy (anonymous):

Yep, you got it. Just to check, your #8 was 5/7 right?

OpenStudy (anonymous):

Yes it was! Thank you sooo much

OpenStudy (anonymous):

Just to check again was #3 b or c?

OpenStudy (anonymous):

#3 was a, when you take the four outside, the x-2 cancels.

OpenStudy (anonymous):

Okay thanks I kinda got confused

OpenStudy (anonymous):

Your welcome!